A and B are two different chemical species undergoing first order decomposition with rate constants ka and kb which are in the ratio 3:2 in magnitude.If the initial concentration of A and B are in the ratio 3:2,what would be the ratio of concentrations of A and B after 3 half-lives of A?

Grade:11

1 Answers

Vikas TU
7 years ago
A’s Half life =  2 years {let}
B’s Half life = 3 years {let} 
After 3 half lives =>
Concentration of A = A/8 and time taken = 2*3 = 6 years.
B will be B/2 at 3 years and after 6 years it would be B/.
Ratio =  (A/8)/(B/4) = > 3/4